OpenStudy (anonymous):

Okay, so I am suppose to solve for x in this problem, but I dont understand the answer. Problem x^2+14x+4=0 Answer -7-3sqrt5, -7-3sqrt5 Can anyone explain?

OpenStudy (vijay):

\[x ={ -b \pm \sqrt{b ^{2} - 4ac}} / 2a \]

OpenStudy (vijay):

b = 14, a = 1 and c = 4, substitute into the above equation, u will fin x.

OpenStudy (anonymous):

\[x^2+14x+4=0\] \[x^2+14x=-4\] \[(x+7)^2=-4+7^2\] \[(x+7)^2=45\] \[x+7=\pm\sqrt{45}\] \[x=-7\pm\sqrt{45}\] and \[\sqrt{45}=\sqrt{9\times 5}=3\sqrt{5}\] so answer is \[x=-7\pm 3\sqrt{5}\]

OpenStudy (anonymous):

you can use quadratic formula too, but you will have a raft of simplification to do at the end

OpenStudy (anonymous):

But how is it separated by a coma? and your answer only has one set of numbers.

OpenStudy (anonymous):

i wrote \[\pm\] meaning both the positive and negative radical

OpenStudy (anonymous):

oh :) Your answer is still different than the answer above

OpenStudy (anonymous):

typo there. it is \[x=-7+3\sqrt{5}\] \[x=-7-3\sqrt{5}\]

OpenStudy (anonymous):

nevermind lol

OpenStudy (anonymous):

that is what the plus/minus indicates. two numbers

OpenStudy (anonymous):

oh. Sorry. It's a new day; therefor I'm basically learning this all over again lol